# Disentangling the Potential Impacts of Papers into Diffusion, Conformity, and Contribution Values

### v3

@@ 5.4.1. Weighting Score Visualization
Figure 5. The outcomes of the linear regression analysis concerning the ”cites”/”is cited by” scores within the field of chemistry. Below the respective sub-figures, we depict the mean Kullback-Leibler (KL) divergences between the target and source distributions, denoted in units of $10^{-4}$.
@@
To effectively visualize the connections between distributions, we gather all scores subsequent to applying softmax based on the target node. These scores are flattened to observe the overarching trend within the latest year. The original attention scores $e$ are averaged across all heads for ease of interpretation. Notably, for the negative effect in the co-citing relationship, we employ reversed values via $1-c$, rendering them positive to reflect the trend. In Figure 5, robustly positive relationships are observed among the sum distribution $\alpha$, original attention distribution $e$, and the distribution of co-cited/citing strengths $c$. It is reasonable to note that the sum distribution closely resembles either of the sourced distributions. When compared to the co-cited/citing distribution, the sum distribution demonstrates greater similarity to the original attention distribution. Intriguingly, within the initial layers of the model, minimal disparities are evident between the original attention scores and the co-cited/citing strengths. This phenomenon is attributed to both the averaging across all heads and the model’s adaptation to the co-cited/citing information during the learning process. Moreover, while the overall trend is evident, a finer examination of the samples reveals that for instances with varying co-cited/citing strengths, their original attention scores tend to be equal, resulting in the displayed horizontal and vertical distribution evident in Figures 5a and 5e. This equal might arise due to the initial state of the attention mechanism in the Graph Attention Network (GAT). However, as the model progresses in depth, noticeable disparities emerge between these distributions in both ”cites” and ”is cited by” edges. In the upper layers, attention scores exhibit a tendency to concentrate on a select few nodes, while in the lower layers, they disperse attention across a wider array of nodes. Co-cited/citing strengths represent 2-hop information conveyed through intermediate nodes, resembling lower-hop information within GNNs. Therefore, continuously incorporating the co-cited/citing distribution across all layers reinforces signals originating from pivotal nodes across a broader spectrum. It can alleviate the over-smoothing issues by preventing the dominance of a select few nodes within aggregated information. This approach serves to explain why a layer count of 4 emerges as optimal—it encapsulates higher-order information beyond the 2-hop range, allowing the co-cited/citing information to complement the original attention scores. This collaboration achieves multi-hop information integration, mitigating both forgetting and overemphasis within the learning process.
5.4.2. Disentanglement Correlation Analysis
Table 6. Results of Disentanglement Correlation in Chemistry
dif con contri dif (%) con (%) contri (%) citations pred real
dif 0.9463 0.9675 0.6177 -0.5847 0.3845 0.2159 0.9947 0.7804
con 0.9463 0.8787 0.5178 -0.3892 0.1889 0.2765 0.9479 0.7501
contri 0.9675 0.8787 0.5182 -0.6342 0.5124 0.1909 0.9828 0.7790
dif (%) 0.6177 0.5178 0.5182 -0.6870 0.2795 0.0651 0.5641 0.3965
con (%) -0.5847 -0.3892 -0.6342 -0.6870 -0.8897 -0.0157 -0.5760 -0.4305
contri (%) 0.3845 0.1889 0.5124 0.2795 -0.8897 -0.0201 0.4067 0.3197
citations 0.2159 0.2765 0.1909 0.0651 -0.0157 -0.0201 0.2226 0.2025
pred 0.9947 0.9479 0.9828 0.5641 -0.5760 0.4067 0.2226 0.7895
real 0.7804 0.7501 0.7790 0.3965 -0.4305 0.3197 0.2025 0.7895
@@ 5.4.2. Disentanglement Correlation Analysis
In Table 6, we depict the correlations among all disentangled perspective values/proportions and the predicted/real values, focusing on samples with all positively predicted values within the Chemistry test set. For papers containing negative disentangled values, they always have nearly no increased citations. Remarkably, over 80% of them feature negative because their contribution values are the lowest across all perspectives, demonstrating the rationality of our proposed method. Among the correlations observed among the values of all three perspectives, ”conformity” stands out as particularly distinct. Its correlation with ”contribution” is notably the lowest, falling below 0.9. Additionally, its correlation with the predicted/real values also demonstrates a significant gap (-0.05/-0.03) when compared to other perspectives. Despite this, it emerges as the most closely associated perspective with accumulated citations, implying its significance in representing information linked to a paper’s reputation. The predicted increment appears as the most influential predictor of the real increment. Its aggregated value outperforms individual perspectives, highlighting the effectiveness of the proposed disentangled learning approach. Moreover, the real increment demonstrates minimal correlation with the paper’s accumulated citations. This revelation underscores two key points: (1) Citation increment proves to be an effective metric in portraying a paper’s potential impact, with accumulated citation accounting for only a fraction of the future increment. (2) The DPPDCC model adeptly captures the trend in a paper’s citation increment, validating its substantial potential and reliability. In our calculations of the proportions of all three perspectives within predicted values, a notable finding emerges: conformity diffusion stands out as the only perspective with numerous negative correlations. This observation suggests that in the field of Chemistry, papers solely characterized by accumulated citations tend to attract fewer new citations. Moreover, it shows an almost negative correlation with the proportion of contribution (-0.89), indicating their mutual exclusivity. Consequently, conformity and contribution have learned entirely distinct facets of citation increments. Overall, the diffusion perspective can be seen as the most pivotal factor for prediction, as it showcases the highest correlations in both values and proportions. This prominence aligns with the intuitive connection between citation increments and the information diffusion process. Conversely, the secondary perspective, contribution, likely delves into factors beyond the graph structure, aiming to uncover additional crucial elements. As a complement to the numerical citation increment, it offers a more objective evaluation, potentially aiding in the identification of valuable papers by encompassing diverse and complementary information.
5.4.3. Disentangled Composition Visualization
(a) time trend
(b) group detail
(c) value bin
@@ 5.4.3. Disentangled Composition Visualization
Figure 6. Visualization of disentangled value proportions in Chemistry. (a) displays the trend that evolved with the publication time. (b) demonstrates the detailed composition of papers categorized into previous, fresh, and immediate ones. (c) is binning the samples based on the predicted values.
We visualize the disentanglement composition with all positive values in the test set of Chemistry. For the stacked plot associated with time in Figure 6a, a clear trend emerges: the contribution proportion gradually increases, while the conformity proportion undergoes a significant drop. This trend is sensible since old papers tend to accumulate substantial citations, thus attracting more researchers to cite them due to inertia. Additionally, the diffusion proportion appears to be more stable across different times, as it represents the potential hidden properties within diffusion. This stability becomes more pronounced when papers are categorized into previous, fresh, and immediate ones in Figure 6b. Diffusion proportions remain steady in both mean and standard deviation. Moreover, contribution emerges as the most critical perspective, with its dominant proportion. It further holds particular significance for fresh and immediate papers. For immediate papers receiving almost no citations, both the diffusion and conformity perspectives struggle to offer predictive insights. However, the contribution enables the model to anticipate the potential of samples in advance by analyzing their citation context. This capability enhances the model’s robustness, particularly for challenging samples close to cold-start scenarios. Furthermore, in Figure 6c, when we bin the samples based on the predicted logarithmic values, we observe a descending trend for conformity and a growing trend for contribution. Notably, within the lowest bin group, the conformity proportion surpasses the contribution proportion for the first time. The alignment between time and impact groups can unveil an overall trend: in dynamic contexts, previous papers tend to attract fewer new citations and may primarily owe to their conformity, whereas recent and immediate papers are more likely to garner increased citations owing to their genuine contributions.
5.4.4. Disentangled Conformity Representation Visualization
(a) total
(b) prev
(c) fresh
(d) immed
Figure 7. Results of T-SNE visualization for representations from conformity perspective in Chemistry. Various colors correspond to the binning labels assigned to the samples.
To validate that disentangled perspectives learn distinct aspects of the features, we visualize the representations extracted from the conformity encoder with T-SNE employed by scikit-learn (Pedregosa et al., 2011), as this perspective is the only one with available labels. Initially, the T-SNE algorithm is executed on the complete test sets of each category. Subsequently, 3000 points are randomly sampled to generate the figures. The points are color-coded based on the binning labels of the cumulative citations at the test time point. Figure 7 illustrates a consistent and structured distribution of points across all figures, indicating our model’s effective capture of group discrepancies across various binning labels in all categories. Across different categories, variations in the distribution of binning labels are evident due to differences in cumulative citations over time. Fresh Papers published within the last five years, particularly immediate ones, tend to fall under the lowest accumulated citation label (red), dominating a significant area. Nevertheless, discernible boundaries persist between distinct binning labels within these categories. This suggests the model’s capacity to generalize and transfer learned group differences to unseen samples, showcasing its potential applicability in real-world scenarios.
5.5. Case Study (RQ4)
@@
242481 2012 28 43 1.2272 0.8977 1.4735 34.10% 24.95% 40.95% 3.5984 3.0445
15676 2001 19 103 1.2151 0.9443 1.4569 33.60% 26.11% 40.29% 3.6163 3.0445
374215 2015 60 26 1.1949 0.8409 1.6318 32.58% 22.93% 44.49% 3.6675 3.2189
(a) 242481-$c$
(b) 242481-$e$
(c) 242481-$\alpha$
(d) 15676-$c$
(e) 15676-$e$
(f) 15676-$\alpha$
(g) 374215-$c$
(h) 374215-$e$
(i) 374215-$\alpha$
@@
Figure 8. Weighted paper ”cites” subgraphs of selected cases organized by co-cited strengths $c$, average original attention scores $e$, and sum scores $\alpha$ in Chemistry. we filter the edges within 2-hop graphs by retaining only the top 30% accumulated weights of each node and preserve the direct citation or reference nodes of the target paper to attain these sparser graphs. We only label the target paper and its directly linked highly-cited citations.
@@
To delve into the intricate workings of our proposed model in addressing RQ4, we select three representative papers for each perspective. These papers should be co-cited or citing papers with similar topics, reflecting comparable predicted and real potential impact. Our approach involves contrasting their metadata, disentangled values, perspective proportions, and predicted/real potential impact. In this section, we visualize graph structures organized by co-cited strengths $c$, average original attention scores $e$, and sum scores $\alpha$ within ”cites” edge subgraphs of the selected papers. To maintain clarity, we focus on graphs extracted from the latest year of the snapshots and the final layer of DPPDCC. We then filter the edges in 2-hop graphs based on the top 30% accumulated weights, resulting in sparser graphs. To create target-centric connected graphs, we retain only the direct citations and references of the target paper. Highly-cited papers, representing papers that provide the top 80% of citations within their respective fields of the whole datasets, are highlighted by filling in red within the graphs. Our labeling strategy encompasses the target paper and its directly connected highly-cited citations.
@@
As outlined in Table 7, 242481, 15676, and 374215 respectively embody the diffusion, conformity, and contribution perspectives. Despite their near-identical predicted and real increments, their disentangled value compositions differ from one another. Notably, 374215 stands as a recent paper not encountered during the training phase, while 242481 and 15673 are earlier publications, each with varying years since publication and distinct accumulated citations. These papers exhibit perspective proportions aligning with all trends discussed in Section 5.4.3. As depicted in Figure 8, $c$ subgraphs manifest as the densest, while $e$ and $\alpha$ subgraphs appear considerably sparser, featuring fewer interconnected nodes. In the higher layers of DPPDCC, the model tends to concentrate on specific nodes, resulting in sparsity in these graphs. The introduced co-cited strengths $c$ serve as a complementary factor to the original attention scores $e$, aiding in identifying valuable citations from highly-cited papers. Notably, $c$ consistently incorporates highly-cited papers overlooked in $e$ subgraphs. For instance, Figures 8b and 8e consider only one highly-cited paper each (532725 and 91549). However, in Figures 8c and 8f, three additional highly-cited papers (559976, 532726, 180237 and 26005, 582958, 193949) - initially presented in Figures 8a and 8d - are taken into account.
@@
For 242481, within the test set, it stands as a recent publication with limited citations due to its relatively short existence. Despite its lower accumulated citations compared to 15676, 242481 has garnered significant attention from four highly-cited papers, equaling the number received by 15676. Consequently, its average influential citation surpasses that of 15676, leading to a higher proportion of the diffusion perspective in its incremental source. Furthermore, since subgraphs depicted in Figure 8 are based on the ”cites” edge, only edges from valuable citations are retained. In contrast to 15676, 242481 boasts numerous connections from references, indicating its significance as an important citation from the perspective of its references. Consequently, it emerges as a pivotal element in the information diffusion process, contributing to its high diffusion values. For 15767, being the oldest paper with a publication span of over 10 years by the test time point (2017), its conformity value stands out as the highest among all cases, contributing significantly to a larger proportion of its incremental impact. Despite its extensive citation count, considering the evolving context of the citation network, its significance has diminished. Over time, citations toward this paper tend to lean more towards its established reputation rather than other factors. For 374215, it is the most recent paper with limited citations. Because it has neither highly-cited citations nor enough accumulated citations, compared to either 242481 or 15767, both its diffusion and conformity perspectives are low in values and proportions. Conversely, its contribution perspective appears as the most pivotal, showcasing the largest values and proportion. Benefiting from its contribution, 374215 successfully surpasses the other two papers in predicted/real potential impact. Interestingly, beyond the factors highlighted in the graph structures or accumulated citation counts, something intrinsic to its contribution stands out, as delineated by DPPDCC. This factor distinguishes itself from mere popularity, emphasizing DPPDCC’s substantial potential for practical application and impact.
@@
Indeed, our proposed DPPDCC effectively disentangles citation increments derived from various factors for the target paper. This disentanglement, reflected in the values and proportions assigned to distinct perspectives, provides more comparable metrics for identifying valuable papers. In future work, we aim to extend this framework and assess its practical application in uncovering novel insights within real-world contexts. This ongoing exploration will enable us to delve deeper into new discoveries and applications facilitated by our disentanglement approach.
6. Conclusion
In this study, we propose a novel model named DPPDCC that disentangles the potential impacts of papers into diffusion, conformity, and contribution values. To better align with real-world conditions, we reformulate the problem by partitioning the datasets based on the observation time point, without excluding lowly-cited papers. Further, DPPDCC comprises two principal components: Citation-aware GNN Encoder (CGE) and Popularity-aware Disentanglement Module (PDM). Given the dynamic heterogeneous graph of a target paper, CGE employs CompGAT to encode the comparative and co-cited/citing information between papers and utilizes Type-specific Snapshot Readout to aggregate the information from each snapshot in an evolutionary manner. Subsequently, PDM identifies amplifying effects in information diffusion and addresses the Matthew effect stemming from accumulated citations from a collective conformity perspective. Finally, it constrains each perspective to address distinct aspects for preserving the genuine contribution. Experimental results on three datasets demonstrate that DPPDCC significantly outperforms alternative approaches for previous, fresh, and immediate papers. Disentanglement visualizations and a detailed case study provide additional evidence supporting DPPDCC’s reliable predictive capabilities.

### v4

@@ G.1. Weighting Score Visualization
Figure 8. The outcomes of the linear regression analysis concerning the ”cites”/”is cited by” scores within the field of computer science. Below the respective sub-figures, we depict the mean Kullback-Leibler (KL) divergences between the target and source distributions, denoted in units of $10^{-4}$.
@@
To effectively visualize the connections between distributions, we gather all scores after applying softmax based on the target node. These scores are flattened to observe the overarching trend within the latest year. The original attention scores $e$ are averaged across all heads for ease of interpretation. Notably, we employ reversed values via $1-c$ for the negative effect in the co-citing relationship, rendering them positive to reflect the trend. In Figure 8, robustly positive relationships are observed among the sum distribution $\alpha$, original attention distribution $e$, and the distribution of co-cited/citing strengths $c$. It is reasonable to note that the sum distribution closely resembles either of the sourced distributions. Compared to the co-cited/citing distribution, the sum distribution demonstrates greater similarity to the original attention distribution. Intriguingly, within the initial layers of the model, minimal disparities are evident between the original attention scores and the co-cited/citing strengths. This phenomenon is attributed to both the averaging across all heads and the model’s adaptation to the co-cited/citing information during the learning process. Moreover, while the overall trend is evident, a finer examination of the samples reveals that for instances with varying co-cited/citing strengths, their original attention scores tend to be equal, resulting in the displayed horizontal and vertical distribution evident in Figures 8a and 8e. This equal might arise due to the initial state of the attention mechanism in the GAT. However, as the model progresses in-depth, noticeable disparities emerge between these distributions in both ”cites” and ”is cited by” edges. In the upper layers, attention scores tend to concentrate on a select few nodes, while in the lower layers, they disperse attention across a wider array of nodes. Furthermore, co-cited/citing strengths represent 2-hop information conveyed through intermediate nodes, resembling lower-hop information within GNNs. Therefore, continuously incorporating the co-cited/citing distribution across all layers reinforces signals from pivotal nodes across a broader spectrum. It can alleviate the over-smoothing issues by preventing the dominance of a select few nodes within aggregated information. This approach serves to explain why a layer count of 4 emerges as optimal—it encapsulates higher-order information beyond the 2-hop range, allowing the co-cited/citing information to complement the original attention scores. This collaboration achieves multi-hop information integration, mitigating both forgetting and overemphasis within the learning process.
G.2. Disentanglement Correlation Analysis
Table 9. Results of Disentanglement Correlation in Computer Science
CS dif con contri dif (%) con (%) contri (%) citations final real
dif 0.9579 0.9990 0.0389 -0.3237 0.4556 0.3209 0.9971 0.8276
con 0.9579 0.9620 -0.1176 -0.1544 0.3495 0.3575 0.9765 0.8091
contri 0.9990 0.9620 0.0276 -0.3236 0.4665 0.3239 0.9981 0.8287
dif (%) 0.0389 -0.1176 0.0276 -0.7657 0.1967 -0.0364 -0.0022 -0.0003
con (%) -0.3237 -0.1544 -0.3236 -0.7657 -0.7812 -0.0056 -0.2861 -0.2477
contri (%) 0.4556 0.3495 0.4665 0.1967 -0.7812 0.0439 0.4382 0.3779
citations 0.3209 0.3575 0.3239 -0.0364 -0.0056 0.0439 0.3331 0.2982
final 0.9971 0.9765 0.9981 -0.0022 -0.2861 0.4382 0.3331 0.8298
real 0.8276 0.8091 0.8287 -0.0003 -0.2477 0.3779 0.2982 0.8298
CHM dif con contri dif (%) con (%) contri (%) citations final real
dif 0.8441 0.9932 0.6134 -0.3982 -0.0478 0.1844 0.9828 0.7834
con 0.8441 0.8723 0.2398 -0.0062 -0.3362 0.2581 0.9267 0.7317
contri 0.9932 0.8723 0.5724 -0.3795 -0.0275 0.2035 0.9911 0.7912
dif (%) 0.6134 0.2398 0.5724 -0.9104 0.4773 -0.0080 0.5075 0.4006
con (%) -0.3982 -0.0062 -0.3795 -0.9104 -0.7981 0.0477 -0.2885 -0.2331
contri (%) -0.0478 -0.3362 -0.0275 0.4773 -0.7981 -0.0897 -0.1263 -0.0883
citations 0.1844 0.2581 0.2035 -0.0080 0.0477 -0.0897 0.2189 0.2022
final 0.9828 0.9267 0.9911 0.5075 -0.2885 -0.1263 0.2189 0.7955
real 0.7834 0.7317 0.7912 0.4006 -0.2331 -0.0883 0.2022 0.7955
PSY dif con contri dif (%) con (%) contri (%) citations final real
dif 0.8642 0.9972 0.5355 -0.5871 0.6302 0.3556 0.9903 0.8196
con 0.8642 0.8740 0.2039 -0.2505 0.2997 0.4998 0.9244 0.7598
contri 0.9972 0.8740 0.5014 -0.5609 0.6149 0.3572 0.9928 0.8199
dif (%) 0.5355 0.2039 0.5014 -0.9869 0.9337 -0.0134 0.4522 0.3917
con (%) -0.5871 -0.2505 -0.5609 -0.9869 -0.9792 0.0056 -0.5070 -0.4328
contri (%) 0.6302 0.2997 0.6149 0.9337 -0.9792 0.0043 0.5570 0.4683
citations 0.3556 0.4998 0.3572 -0.0134 0.0056 0.0043 0.4017 0.3416
final 0.9903 0.9244 0.9928 0.4522 -0.5070 0.5570 0.4017 0.8256
real 0.8196 0.7598 0.8199 0.3917 -0.4328 0.4683 0.3416 0.8256
In Table 9, we illustrate the correlations among all disentangled perspective values/proportions and the predicted/real values, focusing on samples with all positively predicted values across all datasets.
@@ G.2. Disentanglement Correlation Analysis
Taking the correlation analysis of Computer Science as an example, ”conformity” stands out among the values of all three perspectives. Its correlation with the predicted/real values shows a significant gap (-0.02) compared to other perspectives. Despite this, conformity emerges as the most closely associated perspective with accumulated citations, implying its significance in representing information about a paper’s reputation. The predicted increment appears as the most influential predictor of the real increment. Its aggregated value outperforms individual perspectives, highlighting the effectiveness of the proposed disentangled learning approach. Moreover, the real increment demonstrates minimal correlation with the paper’s accumulated citations. This revelation underscores two key points: (1) Citation increment proves to be an effective metric in portraying a paper’s potential impact, with accumulated citation accounting for only a fraction of the future increment. (2) The DPPDCC model adeptly captures the trend in a paper’s citation increment, validating its substantial potential and reliability. In our calculations of the proportions of all three perspectives within predicted values, we find that conformity diffusion is the only perspective with numerous negative correlations. This suggests that in the field of Computer Science, papers characterized solely by accumulated citations tend to attract fewer new citations. Additionally, it shows an almost negative correlation with the proportion of contribution (-0.78), indicating their mutual exclusivity. Consequently, conformity and contribution have learned entirely distinct facets of citation increments. Overall, the contribution perspective is the most pivotal factor for prediction, showcasing the highest correlations in both values and proportions. However, the correlations remain a flexible threshold, where other perspectives also help determine the predicted increments. As a complement to the numerical citation increment, this approach offers a more objective evaluation, potentially aiding in identifying valuable papers by encompassing diverse and complementary information.
In summary, contribution is the most significant perspective across all datasets. Its disentangled value has the largest correlation with both the predicted and real citation increments, as does its proportion in Computer Science and Psychology. Moreover, diffusion and conformity focus on different aspects of the popularity factor. Accumulated citations consistently explain the trend of conformity values, while diffusion complements the structural influence with more flexibility regarding graph features.
G.3. Complete Disentanglement Composition
(a) time trend (CS)
(b) group detail (CS)
(c) value bin (CS)
(d) time trend (CHM)
(e) group detail (CHM)
(f) value bin (CHM)
(g) time trend (PSY)
(h) group detail (PSY)
(i) value bin (PSY)
@@ G.3. Complete Disentanglement Composition
Figure 9. Visualization of disentangled value proportions. (a)/(d)/(g) displays the trend that evolved with the publication time. (b)/(e)/(h) demonstrates the detailed composition of papers categorized into previous, fresh, and immediate ones. (c)/(f)/(i) is binning the samples based on the predicted values.
Figure 9 illustrates the disentangled compositions across all datasets. We observe common trends, such as the similar pattern of conformity and the dominance of contribution. Additionally, details such as the respective slopes and nuanced composition for various categories vary according to the intrinsic characteristics of the corresponding fields.
Appendix H Case Study
@@
981400 2016 28 7 (3/2) 1.2175 0.7428 1.1862 38.69% 23.61% 37.70% 3.1466 3.0445
47429 2009 12 51 (15/3) 1.1432 0.7895 1.1993 36.50% 25.21% 38.29% 3.1320 3.1355
88781 2014 37 5 (1/0) 1.1984 0.7247 1.2307 38.00% 22.98% 39.02% 3.1538 4.7449
(a) 981400-$c$
(b) 981400-$e$
(c) 981400-$\alpha$
(d) 47429-$c$
(e) 47429-$e$
(f) 47429-$\alpha$
(g) 88781-$c$
(h) 88781-$e$
(i) 88781-$\alpha$
@@
Figure 10. Weighted paper ”cites” subgraphs of selected cases organized by co-cited strengths $c$, average original attention scores $e$, and sum scores $\alpha$ in Computer Science. We filter the edges within 2-hop graphs by retaining only the top 30% accumulated weights of each node and preserve the direct citation or reference nodes of the target paper to attain these sparser graphs. We only label the target paper and its directly linked highly-cited citations.
@@
To delve into the intricate workings of our proposed model in addressing RQ4, we select three representative papers for each perspective. These papers should be co-cited or citing papers with similar topics, reflecting comparable predicted and real potential impact. Our approach involves contrasting their metadata, disentangled values, perspective proportions, and predicted/real potential impact. In this section, we visualize graph structures organized by co-cited strengths $c$, average original attention scores $e$, and sum scores $\alpha$ within ”cites” edge subgraphs of the selected papers. To maintain clarity, we focus on graphs extracted from the latest year of the snapshots and average scores across all layers of DPPDCC. We then filter the edges in 2-hop graphs based on the top 30% accumulated weights, resulting in sparser graphs. To create target-centric connected graphs, we retain only the direct citations and references of the target paper. Highly-cited papers, representing papers that provide the top 80% of citations within their respective fields of the whole datasets, are highlighted by filling in red within the graphs. Our labeling strategy encompasses the target paper and its directly connected highly-cited citations.
@@
As outlined in Table 10, papers 981400, 47429, and 88781, respectively, embody the diffusion, conformity, and contribution perspectives. Despite their near-identical predicted increments, their disentangled value compositions differ from one another. Notably, 981400 and 88781 are recent papers not encountered during the training phase, while 47429 is an earlier publication, each with varying years since publication and distinct accumulated citations. These papers exhibit perspective proportions aligning with all trends discussed in Section E. As depicted in Figure 10, $c$ subgraphs manifest as the densest, while $e$ and $\alpha$ subgraphs appear considerably sparser, featuring fewer interconnected nodes. In the higher layers of DPPDCC, the model tends to concentrate on specific nodes, resulting in sparsity in these graphs. The introduced co-cited strengths $c$ act as a complementary factor to the original attention scores $e$, facilitating the identification of significant citations from highly-cited papers. Notably, $c$ consistently integrates significant highly-cited papers that were overlooked, replacing the original set with more valuable ones in $e$ subgraphs. For instance, Figure 10b considers only one highly-cited paper (541312). However, in Figure 10c, one additional highly-cited paper (437859) - initially presented in Figure 10a is taken into account. Similarly, although Figure 10c and 10b share the same number of highly-cited papers (3), the co-cited subgraph $c$ replaces one of them (37213 $\rightarrow$ 51553).
@@
For paper 981400, within the test set, it stands as a recent publication with limited citations due to its relatively short existence. Although it has fewer accumulated citations than paper 47429, 981400 has attracted significant attention from two highly-cited papers, which is noteworthy given its total of seven citations. Consequently, its average influential citation (28.6%) exceeds that of 47429 (5.9%), resulting in a higher proportion of the diffusion perspective in its incremental source. This positions it as a crucial element in the information diffusion process, contributing to its high diffusion values. On the other hand, paper 47429, the oldest paper in the test set with a publication span of nearly 8 years by 2017, stands out with the highest conformity value among all cases, significantly contributing to a larger proportion of its incremental impact. Despite its extensive citation count, its significance has diminished over time in the evolving context of the citation network, with citations tending to emphasize its established reputation rather than other factors. Paper 88781, also recent with limited citations, lacks both highly-cited and accumulated citations compared to papers 981400 or 47429. Consequently, it exhibits low values and proportions in both diffusion and conformity perspectives. However, its contribution perspective emerges as pivotal, showcasing the largest values and proportion. Leveraging its contribution, 88781 surpasses the other two papers in predicted/real potential impact. Interestingly, DPPDCC highlights intrinsic factors to its contribution beyond popularity-related factors highlighted in graph structures or accumulated citation counts, emphasizing DPPDCC’s substantial potential for practical application and impact.
@@
Indeed, our proposed DPPDCC effectively disentangles citation increments from different factors for the target paper. This disentanglement, as evidenced by the values and proportions assigned to distinct perspectives, offers more comparable metrics for identifying valuable papers. In future endeavors, we aspire to expand this framework and evaluate its practical utility in uncovering novel insights within real-world scenarios. This continued exploration will allow us to delve further into discoveries and new applications facilitated by our disentanglement approach.
Appendix I Limitations